# RoundaboutHD: High-Resolution Real-World Urban Environment Benchmark for Multi-Camera Vehicle Tracking

### v1

@@ Abstract
The multi-camera vehicle tracking (MCVT) framework holds significant potential for smart city applications, including anomaly detection, traffic density estimation, and suspect vehicle tracking. However, current publicly available datasets exhibit limitations, such as overly simplistic scenarios, low-resolution footage, and insufficiently diverse conditions, creating a considerable gap between academic research and real-world scenario. To fill this gap, we introduce RoundaboutHD, a comprehensive, high-resolution multi-camera vehicle tracking benchmark dataset specifically designed to represent real-world roundabout scenarios. RoundaboutHD provides a total of 40 minutes of labelled video footage captured by four non-overlapping, high-resolution (4K resolution, 15 fps) cameras. In total, 512 unique vehicle identities are annotated across different camera views, offering rich cross-camera association data. RoundaboutHD offers temporal consistency video footage and enhanced challenges, including increased occlusions and nonlinear movement inside the roundabout. In addition to the full MCVT dataset, several subsets are also available for object detection, single camera tracking, and image-based vehicle re-identification (ReID) tasks. Vehicle model information and camera modelling/ geometry information are also included to support further analysis. We provide baseline results for vehicle detection, single-camera tracking, image-based vehicle re-identification, and multi-camera tracking. The dataset and the evaluation code are publicly available at:https://github.com/siri-rouser/RoundaboutHD.git
@@ 1 Introduction
With the continuous expansion of urbanization, the city traffic management is becoming more and more challenging, which is calling for the development of intelligent transportation systems (ITS) within the broader concept of smart cities. A practical approach to construct real-world ITS is leveraging existing traffic cameras, which capture extensive amounts of useful information in their video footage.
@@
A common application of these video resources is tracking vehicles across multiple camera views, which provides vehicle trajectory information to city management authorities. This task, known as multi-target multi-camera tracking (MTMCT), has garnered considerable research interest in recent years [2]. Generally, this MTMCT problem can be divided into 3 distinct sub-problems, as shown in Fig. 1. (1) Object Detection, which detects and identifies vehicles within the camera scenes; (2) Single-Camera Tracking, which tracks vehicles’ movement across sequential frames within each camera’s field of view; and (3) Multi-Camera Re-identification, which matches and associate’s vehicle identities across separate different camera views. This identification can be performed either at an image-level (image-based re-identification) or at a tracklet-level (video-based re-identification), ultimately reconstructing complete vehicle trajectories.
@@
Each of these sub-problems presents significant challenges in its respective domain, leading to the creation of specialized datasets for object detection, tracking, and re-identification tasks [18, 7, 22]. However, there only have limited number of open-source real-world vehicle-specified MTMCT dataset. To the best of the authors’ knowledge, CityFlow [31] is currently the only open-source dataset of this kind.
@@
The main challenges for making a real-world MCVT datasets are the small inter-class variability and large intra-class variability. Vehicles of the same model typically exhibit very subtle differences, making the re-identification process very challenging even for humans. Conversely, the same vehicle can have dramatic appearance difference due to different lighting condition, camera viewpoint and occlusion, which further complicating the match process. Moreover, the dataset creation is also very time-consuming, requiring the annotation and label the link of hundreds or even thousands of vehicles over extended periods. Considering these challenges, we introduce RoundaboutHD to enrich the available MCVT datasets, the geographic layout is illustrated in Fig. 2. RoundaboutHD comprises 40 minutes of annotated video footage recorded simultaneously from four non-overlapping high-resolution cameras (4K, 15 FPS), with 10 minutes of labeled video per camera. The dataset also introduces increased challenges, including dramatic appearance change due the change of vehicle view angle, frequent occlusions and nonlinear vehicle movements. In addition to the primary dataset designed for evaluating MCVT, we provide subsets for object detection, single camera tracking and image-based ReID. Detailed vehicle attributes (including vehicle type, color, make, and model) with comprehensive camera geometry/calibration information are also provided, enabling diverse research within this domain.
@@
Compared to the widely used CityFlow benchmark, RoundaboutHD focus on a single, continuous scene to ensure high temporal consistency, while maintaining a comparable number of annotated vehicle identities (512 vs. 666). Additionally, it emphasizes real-world representativeness by offering longer, higher-resolution video, and richer contextual annotations—including vehicle attributes and camera calibration information. We hope that the release of this dataset will support and push forward research in ITS, especially for the real-world application of MCVT. The dataset, along with ground-truth annotations for detection, single-camera tracking, image-based ReID, and multi-camera tracking, as well as the associated evaluation tools, are publicly available.
2 Related Work
@@
This section reviews and compares several widely used benchmarks for the ReID task, as summarized in Tab. 1. The table splits existing ReID datasets into image-based and video-based (MTMCT), covering vehicle-specified re-identification scenarios.
2.1 Vehicle Image-Based ReID Dataset
@@
The vehicle images-based ReID datasets [22, 23, 21, 20, 16, 12] are designed specifically for ReID tasks at image level, excluding detection and tracking stages. Among those datasets, VeRI-776 [21] is one of the most commonly used datasets for image-based vehicle ReID. It is an extended version of the original VeRI dataset [22], containing images of 776 vehicles captured by different cameras. VERI-WILD [23] is notable for its large scale, comprising images collected from 174 surveillance cameras over one month in uncontrolled real-world conditions. PKU-VehicleID [20] stands out for its inclusion of detailed vehicle model annotations, offering richer contextual information than many other datasets. The VRIC [16] contributes through its realistic and unconstrained variations in resolution, with all images retrieved from the UA-DETRAC vehicle tracking dataset [34].Vehicle-1M [12] is the largest vehicle ReID dataset published to date, containing nearly 1 million images of 55,527 vehicles across 400 different models. These datasets provide a solid foundation for vehicle ReID research and have helped drive progress toward video-based and multi-camera tracking methods.
2.2 Vehicle Video-Based ReID Dataset
@@
Compared to the image-based ReID task, the video-based ReID task adds the temporal consistency that makes it suitable for MTMCT tasks. While several open-source MTMCT datasets exist for person ReID research (e.g. [28, 40]), to the best of the authors’ knowledge, only two publicly available datasets works on vehicle-specified MTMCT: CityFlow [31] and Synthehicle [15]. CityFlow, released through the CVPR AI City Challenge workshop [26], is the most widely used benchmark in MCVT research. It contains 3.58 hours of video from 46 cameras with 229,680 bounding boxes of 666 vehicle identities annotated. However, each camera covers only a short duration, and the video quality is relatively low (960p), which limits its effectiveness in representing real-world urban environments. Synthehicle, on the other hand, is the first synthetic dataset that eliminate the need for manual labelling. Compared to other datasets, it offers a large-scale material with rich annotations, including 2D bounding boxes, 3D positions, segmentation masks, and depth information. Nevertheless, all video content is generated in the CARLA [9], introducing a domain gap that reduces its applicability to real-world scenarios. In addition to these open-source datasets, several private video-based ReID datasets exist for MCVT evaluation, such as the Large-scale High-resolution Traffic Video (LHTV) dataset [35] and the Highway Surveillance Traffic (HST) dataset [38]. However, these datasets are not publicly available, which restricts their use in broader academic research and benchmarking.
3 Dataset Analysis
@@ 3.1 Dataset Overview
The proposed RoundaboutHD dataset was collected in a small town in the United States. The geographic layout, inter-camera distances, and sample views from each camera are illustrated in Fig. 2. The maximum distance between two cameras is approximately 0.92 km. Several retail parks and side roads are between camera pairs, increasing the uncertainty of vehicle movement patterns and making the tracking environment closer to real-world urban traffic scenarios.
The total duration of the RoundaboutHD dataset is 40 minutes, each 4k 15fps camera contains 10 minutes videos. All videos begin recording simultaneously to support temporally aligned multi-camera tracking. Within the 40-minute recording, a total of 549,909 bounding boxes of 512 vehicles are manually annotated. The number of annotated vehicles per camera is summarized in Tab. 2.
@@
Table 2: Number of Annotated Vehicles in Each Camera. Multi-Cam refers to vehicles appearing in at least two cameras.
Camera modelling parameters are provided based on the camera model in [10], through those parameters, user can transform point coordinates in the image to real-world or to geographic coordinates. An illustration of this mapping process is presented in Fig. 3. In addition, human annotators manually label vehicle attributes including colour, type, make, and specific model for all vehicles shown up in our dataset. This rich contextual information offers valuable support for a range of future research. Based on these annotations, the five most common vehicle makes are Chevrolet, Ford, Honda, Toyota, and Nissan, while the most frequently observed vehicle models are Chevrolet Silverado, Ford Transit, Honda CR-V, Chevrolet Equinox, and Toyota Camry. A statistical summary of vehicle types and colours is provided in Fig. 4.
@@
For privacy concerns, a specific angle of the cameras is pre-set to make sure no recognizable human face or readable license plate is recorded. A comparative overview of RoundaboutHD and existing datasets is provided in Tab. 1.
@@ 3.2 Dataset Annotation
Manual frame-level annotation for MCVT is highly time-consuming and labor-intensive. To improve efficiency, we adopt a semi-automatic annotation pipeline combining detection, feature extraction, and trajectory generation with manual refinement. First, the YOLOv12x [32] object detector pre-trained on COCO dataset [18] is employed to generate initial bounding boxes. A confidence threshold of 0.35 is applied, and detections with bounding box areas smaller than 900 pixels are filtered out to reduce noise from small or distant objects. All initial detection results are manually reviewed to correct false positives and false negative identities. Next, deep appearance features are extracted using a ResNet-101 [13] model. These features, along with the refined detections, are input to SMILEtrack [33] to get initial trajectories. The resulting trajectories are manually inspected, and ID switches are corrected to ensure accurate single-camera tracking. Based on above information, the cross-camera association is conducted manually. Vehicles are matched across different camera views based on a combination of visual attributes (colour, make, and model) and spatial-temporal consistency between camera pairs. A challenging cross-camera vehicle movement example is shown in Fig. 5. Finally, an image-based ReID subset is constructed based on the multi-camera association result. To ensure high-quality visual data, cropped images are extracted from bounding boxes with a minimum resolution of 3,500 pixels, sampled every 5 frames, to form the training and gallery sets. The query set is sampled every 45 frames with prioritizing high resolution cropped images. This approach ensures sufficient visual diversity and quality for robust vehicle ReID benchmarking.
Figure 5: Example of cross-camera vehicle trajectory annotation with timestamp information.
@@ 3.3 Subset Overview
Single-camera tracking dataset. The SCT subset includes a total of 1,082 annotated trajectories, composed of 233,189 bounding boxes. While Table 2 provides the number of unique vehicle identities per camera, the total number of annotated trajectories slightly exceeds the number of unique identities. The difference is raised from occasional long-term occlusions, which leads to id switch within individual camera views. The distribution of annotated trajectories across the four cameras is summarized in Tab. 3:
@@
3.4 Evaluation
To facilitate standardized benchmarking, we provide an evaluation framework for the main dataset and all subsets as part of the package. This benchmark enables comparison of different algorithm by evaluating the result text file with ground truth using task-specific performance metrics. Details on the required input file formats for each task are available in our open-source GitHub repository.
Object detection evaluation. For object detection, we adopt standard evaluation metrics, including AP@0.5, AP@0.75, AP@0.9, and the mean Average Precision (mAP).
Single-camera tracking evaluation. For SCT, we select serval metrics used by the famous MOT Challenges [7], includes IDF1 score [29], ID Switches and MOTA [3] to evaluate the performance of each SCT algorithm.
@@ 3.4 Evaluation
Overall, this evaluation framework provides a unified, easy-to-use approach for benchmarking algorithms across all tasks on the RoundaboutHD dataset.
@@ 4.1 Object Detection
Figure 6: Example Challenging Cases for Object Detection in RoundaboutHD. (a) A combination of rare vehicle types and severe occlusion leads to both false positives and false negatives. (b) Special vehicle appearances result in incomplete detections. (c) Heavy occlusion causes false negatives.
Object detectors are typically classified into two categories: one-stage detectors, such as YOLO [32], and two-stage detectors, such as Fast R-CNN [11]. Since 2020, one-stage detectors—particularly the YOLO family—have become dominant in both academic research and real-world applications due to their speed and accuracy. To evaluate how SOTA detectors perform under such conditions, we use the recently released attention-based YOLOv12x [32] model and transformer architecture based YOLOv11X [17] as a benchmark, setting confidence threshold to 0.3, non-maximum suppression (NMS) to be true, using an inference image size of 1280, and applying an IoU threshold of 0.4. Tab. 4 and Tab. 5 presents its performance on the RoundaboutHD dataset, offering insights into detection performance in complex urban scenes.
@@
According to the results, the transformer-based YOLOv11 slightly outperforms YOLOv12, achieving a 1.38% higher average mAP. Notably, the ground truth annotations exclude static vehicles parked in the visible parking lot. However, the YOLO detector still identifies these vehicles, leading to a higher FP rate—particularly in Cam01, where multiple vehicles on the right side of the video footage are obvious. This affects the apparent detection performance and should be considered when interpreting the results.
Among the four cameras, Cam04 is the least affected by this issue and thus provides a more representative measure of detection accuracy. Incorporating a filtering mechanism to remove static vehicle detections should be considered for future experiments to obtain more reliable and meaningful evaluation results.

### v2

@@ Abstract
The multi-camera vehicle tracking (MCVT) algorithms hold significant potential for various smart city applications, including anomaly detection, traffic density estimation, and suspect vehicle tracking. Developing and benchmarking such algorithms needs high quality opensource datasets. However, the publicly available datasets at the moment exhibit several limitations, such as overly simplistic scenarios, low-resolution footage, and insufficiently diverse conditions, creating a considerable gap between academic research and real-world scenarios. To fill this gap, we introduce RoundaboutHD, a comprehensive, high-resolution multi-camera vehicle tracking benchmark dataset specifically designed to represent real-world roundabout scenarios. RoundaboutHD provides a total of 40 minutes of labeled video footage captured by four non-overlapping, high-resolution (4K resolution, 15 fps) cameras. In total, 512 unique vehicle identities are annotated across different camera views, offering rich cross-camera association data. In addition to the full MCVT dataset, several subsets are also available for object detection, single camera tracking, and image-based vehicle re-identification (ReID) tasks. Vehicle model information and camera modeling/geometry information are also included to support further analysis. We provide baseline results for vehicle detection, single-camera tracking, image-based vehicle re-identification, and multi-camera tracking. The dataset and the evaluation code are publicly available at: https://github.com/siri-rouser/RoundaboutHD.git
@@ 1 Introduction
With the continuous expansion of urbanization, the city traffic management is becoming more and more challenging, motivating the development of intelligent transportation systems (ITS) within the broader concept of smart cities. A practical approach to construct ITS is taking advantage of existing traffic cameras, which capture extensive amounts of useful information from video footage.
@@
A common application of these video resources is tracking vehicles across multiple camera views, which provides vehicle trajectory information to city management authorities. This task, known as multi-target multi-camera tracking (MTMCT), has attracted considerable research interest in recent years [2]. Generally, this MTMCT problem can be divided into 3 distinct sub-problems, as shown in Fig. 1. (1) Object Detection, which detects and identifies vehicles within the camera scenes; (2) Single-Camera Tracking, which tracks vehicles’ movement across sequential frames within each camera’s field of view; and (3) Multi-Camera Re-identification, which matches and associate’s vehicle identities across separate different camera views. This identification can be performed either at an image-level (image-based re-identification) or at a tracklet-level (video-based re-identification), ultimately reconstructing complete vehicle trajectories.
@@
Each of these sub-problems presents significant challenges in its respective domain, leading to the creation of specialized datasets for object detection, tracking, and re-identification tasks [18, 7, 22]. However, there is a very limited number of open-source real-world vehicle-specified MTMCT dataset has been proposed, and to the best of the authors’ knowledge, CityFlow [31] is currently the only open-source dataset of this kind.
@@
The main challenges in developing a real-world MCVT datasets arise from the small inter-class variability and large intra-class variability. Vehicles of the same model typically exhibit very subtle differences, making the re-identification process very challenging even for humans. Conversely, the same vehicle can have dramatic appearance difference due to different lighting conditions, camera viewpoint and occlusion, which further complicating the match process. Moreover, the dataset creation is also very time-consuming, requiring the annotating, labeling and linking of hundreds or even thousands of vehicles over extended periods. Considering these challenges and to enrich the available MCVT datasets, we introduce RoundaboutHD, a high resolution and challenging dataset comprising of sequences captured from four location in a small town located in US, the geographic layout is illustrated in Fig. 2. RoundaboutHD comprises in total 40 minutes of annotated video footage recorded simultaneously from four non-overlapping high-resolution cameras (4K, 15 FPS), with 10 minutes of labeled video from each camera. The dataset also introduces increased challenges, including dramatic appearance change due the change of vehicle view angle, frequent occlusions and nonlinear vehicle movements. In addition to the primary dataset designed for evaluating MCVT, we provide subsets for object detection, single camera tracking and image-based ReID. Detailed vehicle attributes (including vehicle type, color, make, and model) with comprehensive camera geometry/calibration information are also provided, enabling diverse research within this domain.
@@
Compared to the widely used CityFlow benchmark, RoundaboutHD focus on a single scene with longer duration footage for each camera, while still maintaining a comparable number of annotated vehicle identities (512 vs. 666), demonstrating the substantial scale of the dataset. Additionally, it emphasizes real-world representativeness by offering longer, higher-resolution video, and richer contextual annotations—including vehicle attributes and camera calibration information. We hope that the release of this dataset will support and push forward research in ITS, especially for the real-world application of MCVT. The dataset, along with ground-truth annotations for detection, single-camera tracking, image-based ReID, and multi-camera tracking, as well as the associated evaluation tools, are publicly available.
2 Vehicle ReID Datasets
@@
Tab. 1 compares nine of the most widely used datasts for vehicle ReID. The table splits existing ReID datasets into image-based and video-based (MTMCT), covering vehicle-specified re-identification scenarios, and these scenarios are discussed in more details below.
2.1 Image-Based Vehicle ReID Dataset
@@
The images-based vehicle ReID datasets [22, 23, 21, 20, 16, 12] are designed specifically for ReID tasks at image level, excluding the detection and tracking stage. Among these datasets, VeRI-776 [21] is one of the most widely used benchmarks for image-based vehicle ReID. It extends the original VeRI dataset [22], increasing the number of annotated vehicle identities from 619 to 776, with images captured from 20 camera views. A larger dataset, VERI-WILD [23] is notable for its large scale, comprising images collected from 174 surveillance cameras over one month in uncontrolled real-world conditions. On the other hand, the PKU-VehicleID [20] stands out for its inclusion of detailed vehicle make/model annotations, offering richer vehicle identity information than many other datasets. The VRIC dataset [16] is built from the well-known UA-DETRAC vehicle tracking dataset [34], and aims to reflect more realistic conditions for vehicle ReID. It contains lower-resolution images from multiple viewpoints, lighting conditions and occlusion conditions, which better represents real-world vehicle appearances in wide-area traffic scenes. And finally, Vehicle-1M [12] is the largest vehicle ReID dataset published to date, it contains nearly 1 million images of 55,527 vehicles across 400 different models. All dataset mentioned above provide a solid foundation for image-based vehicle ReID research and have helped to drive the research progress toward the video-based ReID task.
2.2 Video-Based Vehicle ReID Dataset
@@
Compared to image-based ReID, video-based ReID incorporates temporal information and make it corresponds to the MTMCT task. While several MTMCT datasets have been developed for person ReID research (e.g., [28, 40]), vehicle-specified MTMCT dataset remain scarce. To the best of the authors’ knowledge, there are only two publicly available datasets work on vehicle-specified MTMCT: the CityFlow [31] and Synthehicle [15]. The CityFlow dataset, released through the CVPR AI City Challenge workshop [26], has been the most widely used dataset in MCVT research. It contains 3.58 hours of video from 46 cameras with 229,680 bounding boxes of 666 vehicle identities annotated. However, each camera captures only a short duration of footage, and the video quality is relatively low (960p). Given the increasing availability of high-resolution traffic cameras, these limitations reduce CityFlow’s effectiveness in representing modern real-world urban environments. The Synthehicle dataset, on the other hand, is the first synthetic dataset with identities programatically created in CARLA [9] simulation environment which eliminates the need for manual labeling. Compared to other datasets, it offers a large amount of video content with abundant and accurate annotations, including 2D bounding boxes, 3D positions, segmentation masks, and depth information. Nevertheless, as all video content is generated in the virtual environment, this introduces a domain gap that reduces its applicability to real-world scenarios. In addition to these open-source datasets, several private video-based ReID datasets exist for MCVT evaluation, such as the Large-scale High-resolution Traffic Video (LHTV) dataset [35] and the Highway Surveillance Traffic (HST) dataset [38]. Since these datasets are not publicly available, their use is high restricted in broader academic research and benchmarking.
3 The Proposed RoundaboutHD Dataset
@@ 3.1 Dataset Overview
The proposed RoundaboutHD dataset was collected in Carmel, Indiana, a small town located in the United States with a high number of roundabouts not commonly seen in the country. The geographic layout, inter-camera distances, and sample views from each camera are illustrated in Fig. 2. The maximum distance between two cameras is approximately 0.92 km. There are several retail parks and side roads between pairs of cameras, increasing the uncertainty of vehicle movement patterns and making the tracking environment closer to real-world urban traffic scenarios.
@@
Table 2: Number of Annotated Vehicles in Each Camera. Multi-Cam refers to vehicles appeared in at least two cameras.
The total duration of the RoundaboutHD dataset is 40 minutes, each 4k 15fps camera contains 10 minutes videos. All videos begin recording simultaneously to support temporally aligned multi-camera tracking. Within the 40 minutes recording, a total of 549,909 bounding boxes of 512 vehicles are manually annotated. The number of annotated vehicles per camera is summarized in Tab. 2.
Camera parameters are provided based on the model built in [10], through these parameters, users can transform pixel in the image space to 3D space coordinates or to real-world GPS coordinates. An illustration of this mapping process is presented in Fig. 3. In addition, human annotators manually label vehicle attributes including colour, type, make, and specific model for all vehicles identified in the dataset. This rich contextual information offers valuable support for a range of future research. Based on these annotations, the five most common vehicle makes are Chevrolet, Ford, Honda, Toyota, and Nissan, while the most frequently observed vehicle models are Chevrolet Silverado, Ford Transit, Honda CR-V, Chevrolet Equinox, and Toyota Camry. A statistical summary of vehicle types and colours is provided in Fig. 4.
@@
To address the privacy concerns, a specific angle of the cameras is pre-set that ensures no recognizable human face or readable license plate is recorded.
@@ 3.2 Dataset Annotation
Manual frame-level annotation for MCVT is highly time-consuming and labor-intensive. To speed up the process, we adopt a semi-automatic annotation pipeline combining detection, feature extraction, and trajectory generation with manual refinement. First, the YOLOv12x [32] object detector pre-trained on COCO dataset [18] is applied to generate initial bounding boxes. A confidence threshold of 0.35 is applied, and detections with bounding box areas smaller than 900 pixels are filtered out to reduce noise from small or distant objects. All initial detection results are manually reviewed to correct false positives and false negative identities. Next, deep appearance features are extracted using a ResNet-101 [13] model. These features, along with the refined detections, are inputted to SMILEtrack [33] to get initial trajectories. The resulting trajectories are manually inspected and corrected to make the single-camera tracking ground truth. Based on above information, the cross-camera association is conducted manually. Vehicles are matched across different camera views based on a combination of visual attributes (color, make, and model) and spatial-temporal consistency between camera pairs. A challenging cross-camera vehicle movement example is shown in Fig. 5. Finally, an image-based ReID subset is constructed based on the multi-camera association result. To ensure high-quality visual data, cropped images are extracted from bounding boxes with a minimum resolution of 3,500 pixels, sampled every 5 frames, to form the training and gallery sets. The query set is sampled every 45 frames with prioritizing high resolution cropped images. This approach ensures sufficient visual diversity and quality for robust vehicle ReID benchmarking.
Figure 5: A Challenging Example of Cross-Camera Vehicle Trajectory Annotation. In this case, the target vehicle exits the FOV of Cam01 at timestamp 2955 and appears to head toward Cam02. However, it unexpectedly reappears in Cam01 at timestamp 5730, indicating a U-turn. This movement pattern will pose a significant challenge to MCVT algorithms that rely heavily on spatiotemporal information to constrain the search space.
@@ 3.3 Subset Overview
Single-camera tracking dataset. The SCT subset includes a total of 1,042 annotated trajectories, composed of 232,945 bounding boxes. While Tab. 2 provides the number of unique vehicle identities per camera, the total number of annotated trajectories slightly exceeds the number of unique identities. The difference is raised from occasional long-term occlusions—often caused by large vehicles overlapping smaller ones—which leads to id switch within individual camera views. The distribution of annotated trajectories for each of the camera is summarized in Tab. 3.
@@
3.4 Benchmark Evaluation of RoundaboutHD
To facilitate standardized benchmarking, we provide an evaluation framework for the main dataset and all subsets as part of the package. This benchmark enables comparison of different algorithms by evaluating the result text file with ground truth under task-specific performance metrics. Details on the required input file formats for each task are available in our open-source GitHub repository.
Object detection evaluation. For object detection, we adopt standard evaluation metrics based on Average Precision (AP), including AP@0.5, AP@0.75, AP@0.9, and the mean AP.
Single-camera tracking evaluation. For SCT, we select serval metrics used by the famous MOT Challenges [7], including IDF1 score [29], ID Switches and MOTA [3] to evaluate the performance of each SCT algorithm.
@@ 3.4 Benchmark Evaluation of RoundaboutHD
In the next section, the framework will be used in benchmarking experiments that helps analyzing the attributes of RoundaboutHD in comparison with existing datasets.
@@ 4.1 Object Detection
Figure 6: Example Challenging Cases for Object Detection in RoundaboutHD. (a) A combination of rare vehicle types and severe occlusion leads to both false positives and false negatives. (b) Special vehicle appearances result in incomplete detections. (c) Heavy occlusion causes false negative detection.
Object detectors are typically classified into two categories: one-stage detectors, such as YOLO [32], and two-stage detectors, such as Fast R-CNN [11]. Since 2020, one-stage detectors—particularly the YOLO family—have become dominant in both academic research and real-world applications due to their speed and accuracy. To evaluate how SOTA detectors perform in RoundaboutHD, we use the recently released attention-based YOLOv12x [32] model and transformer architecture based YOLOv11X [17] as baseline model, setting the confidence threshold to 0.3, non-maximum suppression (NMS) to be true, using an inference image size of 1280, and applying an IoU threshold of 0.4. Tab. 4 and Tab. 5 presents its performance on the RoundaboutHD dataset, offering insights into detection performance in complex urban scenes.
@@
According to the results, the transformer-based YOLOv11x model slightly outperforms YOLOv12x model, achieving a 1.38% higher average mAP. It is important to note that the ground truth annotations intentionally exclude static vehicles parked in the visible parking lot. However, the YOLO detector still detects these vehicles, which increases the number of false positives. This leads to an apparent drop in detection accuracy and should be considered when interpreting the results.